OpenStudy (anonymous):

Find the radius of a circle if a central angle of 3pi/2 intercepts an arc of 26.4 centimeters. 0.19 cm 3.14 cm 5.6 cm

OpenStudy (jdoe0001):

\(\bf \textit{arc's length}=s=\cfrac{{\color{blue}{ r}\pi \theta}}{180}\implies \cfrac{180s}{\pi \theta}={\color{blue}{ r}}\)

OpenStudy (jdoe0001):

you're given the arc's length and the angle plug them in

OpenStudy (jdoe0001):

hmmm actually your angle is in radians... ok... so one sec

OpenStudy (jdoe0001):

\(\bf \textit{arc's length}=s={\color{blue}{ r}}\theta\implies \cfrac{s}{\theta}={\color{blue}{ r}}\)

OpenStudy (anonymous):

okay thank you!